In Problem 26, your best bet is to sketch a graph on paper with all the known information – you know the coordinates of the solution, and should be able to easily determine the y-intercept of the second equation - important since this would give you two points on the same line, from which you should then be able to find m. In the first equation, you’ll have to do some math to find the value of its slope – but watch out for the value of b, because here (HINT), it may not be the same as the value of the y-intercept. Also, we’re only looking for the values of m and b – no graph needs to be submitted for this problem!
26) Find values for m and b in the following system so that the solution to the system is (-3, 4).
5x + 7y = b
mx + y = 22

Answer by scott8148(6628)

solution is (-3,4) so 5x+7y=b ... 5(-3)+7(4)=b ... -15+28=b ... 13=b

mx+y=22 ... m(-3)+4=22 ... -3m=18 ... m=-6